The sum of the ages of a gold coin and a silver coin is 65 years. The age of the gold coin 10 years from now will be 10 years less than the age of the silver coin 5 years ago. Find the present ages of the two coins. Gold coin --- yr. Silver coin--- yr. I got 1996 for the gold coin and 1971 for the silver coin.
Answer by ankor@dixie-net.com(22740)

The sum of the ages of a gold coin and a silver coin is 65 years.
The age of the gold coin 10 years from now will be 10 years less than the age of the silver coin 5 years ago.
Find the present ages of the two coins. Gold coin --- yr. Silver coin--- yr.
:
Let g = age of the gold coin
Let s = age of the silver
:
Write an equation for each statement:
:
"The sum of the ages of a gold coin and a silver coin is 65 years."
g + s = 65
:
"The age of the gold coin 10 years from now will be 10 years less than the age of the silver coin 5 years ago."
g + 10 = (s-5) - 10
g + 10 = s - 15
g = s - 15 - 10
g = s - 25
Put equation into the standard form, add to the 1st equation
g - s = -25
g + s = 65
---------------addition eliminates s, find g
2g = 40
g = 20 yrs, age of the gold
then
20 + s = 65
s = 65 - 20
s = 45 yrs age of the silver
;
:
See if that works in the statement
" The age of the gold coin 10 years from now will be 10 years less than the age of the silver coin 5 years ago. "
20 + 10 = (45 - 5) - 10
30 = 40 - 10
: